# Contestable AI needs Computational Argumentation

### v1

@@ 4 The Role of Computational Argumentation
We argue that computational argumentation (CA), broadly understood as in (?; ?), is ideally suited to support AI contestability computationally. More strongly, we argue that it is ideally and uniquely suited, as it can cover all aspects of the abstract view in Figure 1 organically. We support this view with reference to several lines of work in the CA literature, as follows.
@@
CA for Explanation. CA has been widely used for XAI (see (?; ?) for recent overviews). It can provide abstractions of several existing, widely used models, e.g. as in (?; ?; ?; ?; ?), and can itself directly serve as the basis of models (?; ?; ?). Natural forms of explanations can be obtained from CA abstractions, e.g. dispute trees (?; ?), defence sets (?) and attribution scores, e.g. gradient-based argument attributions (?), Shapley-based relation attributions (?; ?), amongst several.
@@
CA for Redress. CA provides methods for revising (?; ?; ?) and repairing knowledge bases (?). Much research has been investigated in the context of forgetting (?; ?; ?) and enforcement (?; ?; ?); researchers investigated the effect of expansions (?; ?; ?) and changes in the knowledge base (?; ?; ?; ?). In addition, incomplete argumentation frameworks (?; ?; ?) incorporate uncertainty which enables redress on a conceptual level.
@@
These generic lines of work in CA are useful starting points for supporting redress of models which can be abstracted argumentatively. In addition, reasoning with argumentation frameworks drawn from models is the basis for forms of redress of some (machine learning) models, e.g. as in (?) and (?) for natural language processing, and of some (scheduling) models, as in (?).
@@
CA for Interaction. Interactions between agents, often modelled as dialogues, have been shown to be effectively supported by various forms of CA, e.g. argument schemes (?) or abstract argumentation frameworks (?), in a number of settings. CA’s formal nature allows for principled desiderata of argumentative agent protocols in these settings, as defined in (?). These settings include: computational persuasion (?; ?; ?; ?), framed as selecting the most effective arguments for changing the mind of the other agents; information-seeking and inquiry (?; ?), where agents share information comprising arguments which are private or public; and other areas such as the handling of maliciousness amongst agents (?). Approaches to multi-agent argumentation such as these have been been shown to be useful in various real world applications, e.g. regulatory compliance (?), recommender systems (?; ?; ?; ?) and, more recently, interactive XAI (?; ?; ?). One such approach which looks to have particular promise is that of (?), where argumentative exchanges frame interactive explanation between agents as a conflict resolution problem, while accounting for humans’ cognitive biases. These CA-based approaches to modelling interactions look to have promise for human interaction with Large Language Models, as in (?).
@@
CA and the viewpoint of the contester. CA-based dialogue and interaction methods, e.g. as in the aforementioned (?), already have the potential to accommodate the contester’s viewpoint, leveraging on the use of CA to provide abstractions for underpinning . Indeed, e.g. in (?), contested and contester are interchangeably seen as argumentation frameworks.
5 Conclusion
@@
Our analysis is restricted, for simplicity and brevity, but could be naturally extended to cover broader scenarios. For example, we have focused on two-parties scenarios where a single ADS is contested and another ADS or a human is doing the contesting; however multi-party scenarios are also possible, e.g. when, in addition to a bank customer, a regulator contests a financial institution. Also, in line with the existing literature on contestability, we have assumed that only ADSs can be contested, but it may be useful to consider the possibility that humans may be contested by ADSs, e.g. in the spirit of (?). In addition to broadening our view to these and other more complex scenarios, future work will be needed to provide evidence of our claim, by building forms of contestable AI that naturally use CA.
Finally, we have advocated CA for contestability, but other KR techniques could provide useful support for some contestability aspects. Specifically, formal verification of AI models could be used to ensure the functional correctness of ADSs, e.g. as in (?; ?). Similarly, rigorous logic-based explainability techniques (?; ?; ?) could be used to ensure the faithfulness and trustworthiness of explanations.

### v2

@@ 4 The Role of Computational Argumentation
In this position paper we argue that computational argumentation (CA) is ideally suited to support AI contestability computationally. CA, broadly understood as in (Atkinson et al., 2017; Baroni et al., 2018), is a branch of Knowledge Representation & Reasoning which represents information in terms of arguments and dialectical relations (of attack and, possibly, support) between them. CA is equipped with semantics to reach some form of consensus regarding conclusions to be drawn. As such, CA is ideally and uniquely suited to cover all aspects of the abstract view in Figure 1 organically. We support this view with reference to several lines of work in the CA literature, as follows.
@@
CA for explanation. CA has been widely used for XAI (see (Vassiliades, Bassiliades, and Patkos, 2021; Cyras et al., 2021) for recent overviews). It can provide abstractions of several existing, widely used models, e.g. as in (Potyka, Yin, and Toni, 2023; Ayoobi, Potyka, and Toni, 2023; Prakken and others, 2020; Fan, 2018; Cyras et al., 2019b), and can itself directly serve as the basis of models (Rago, Cocarascu, and Toni, 2018; Rago et al., 2020; Cocarascu et al., 2020). Natural forms of explanations can be obtained from CA abstractions, e.g. dispute trees (Fan and Toni, 2015b; Čyras et al., 2019a), defence sets (Arioua, Tamani, and Croitoru, 2015) and attribution scores, e.g. gradient-based argument attributions (Yin, Potyka, and Toni, 2023), Shapley-based relation attributions (Amgoud, Ben-Naim, and Vesic, 2017; Yin, Potyka, and Toni, 2024), amongst several.
@@
CA for redress. CA provides methods for revising (Snaith and Reed, 2017; Baumann and Brewka, 2015; Falappa, Kern-Isberner, and Simari, 2009) and repairing knowledge bases (Ulbricht and Baumann, 2019). Much research has been investigated in the context of forgetting (Berthold, Rapberger, and Ulbricht, 2023; Baumann, Gabbay, and Rodrigues, 2020) and enforcement (Rapberger and Ulbricht, 2023; Baumann et al., 2021); researchers investigated the effect of expansions (Prakken, 2023; Oikarinen and Woltran, 2011; Cayrol, de Saint-Cyr, and Lagasquie-Schiex, 2010) and changes in the knowledge base (Doutre and Mailly, 2018; Booth et al., 2013; Moguillansky et al., 2013; Niskanen, 2020). In addition, incomplete argumentation frameworks (Baumeister et al., 2021; Alfano et al., 2023) incorporate uncertainty which enables redress on a conceptual level.
@@
These generic lines of work in CA are useful starting points for supporting redress of models which can be abstracted argumentatively. In addition, reasoning with argumentation frameworks drawn from models is the basis for forms of redress of some (machine learning) models, e.g. as in (Carstens and Toni, 2017) and (Freedman et al., 2024) for natural language processing, and of some (scheduling) models, as in (Cyras et al., 2019b).
@@
CA for interaction. Interactions between agents, often modelled as dialogues, have been shown to be effectively supported by various forms of CA, e.g. argument schemes (Panisson, McBurney, and Bordini, 2021) or abstract argumentation frameworks (de Tarlé, Bonzon, and Maudet, 2022), in a number of settings. CA’s formal nature allows for principled desiderata of argumentative agent protocols in these settings, as defined in (McBurney, Parsons, and Wooldridge, 2002). These settings include: computational persuasion (Fan and Toni, 2012; Hunter, 2018; Calegari, Riveret, and Sartor, 2021; Donadello et al., 2022), framed as selecting the most effective arguments for changing the mind of the other agents; information-seeking and inquiry (Black and Hunter, 2007; Fan and Toni, 2015a), where agents share information comprising arguments which are private or public; and other areas such as the handling of maliciousness amongst agents (Kontarinis and Toni, 2015). Approaches to multi-agent argumentation such as these have been been shown to be useful in various real-world applications, e.g. regulatory compliance (Raymond et al., 2022), recommender systems (Briguez et al., 2014; Teze, Godo, and Simari, 2018; Rago et al., 2020, 2021) and, more recently, interactive XAI (Madumal et al., 2019; Calegari et al., 2022; Paulino-Passos and Toni, 2022). One such approach which looks to have particular promise is that of Rago, Li, and Toni (2023), where argumentative exchanges frame interactive explanation between agents as a conflict resolution problem, while accounting for humans’ cognitive biases. These CA-based approaches to modelling interactions look to have promise for human interaction with Large Language Models, as in (Freedman et al., 2024).
@@
CA and the viewpoint of the contester. CA-based dialogue and interaction methods, e.g. as in the aforementioned (Rago, Li, and Toni, 2023), already have the potential to accommodate the contester’s viewpoint, leveraging on the use of CA to provide abstractions for underpinning . Indeed, e.g. in (Rago, Li, and Toni, 2023), contested and contester are interchangeably seen as argumentation frameworks.
5 Discussion and Future Work
@@
Our analysis is restricted, for simplicity and brevity, but could be naturally extended to cover broader scenarios. For example, we have focused on two-party scenarios where a single ADS is contested and another ADS or a human is doing the contesting; however multi-party scenarios are also possible, e.g. when, in addition to a bank customer, a regulator contests a financial institution. Also, in line with the existing literature on contestability, we have assumed that only ADSs can be contested, but it may be useful to consider the possibility that humans may be contested by ADSs, e.g. in the spirit of (Miller, 2023). In addition to broadening our view to these and other more complex scenarios, future work will be needed to provide evidence of our claim, by building forms of contestable AI that naturally use CA. This involves connecting what we see are the main requirements of contestability (explanations, grounds for contestation, redress, and interaction) in concrete applications and scenarios. CA has already been used to solve several individual challenges, e.g. see (Borg and Bex, 2020; Lawrence, Visser, and Reed, 2023; Cyras et al., 2019b). However, considerable engineering work is needed to combine solutions towards fully-fledged end-to-end contestable systems. For instance, in order to support contestability with modern AI systems, we envisage that considerable effort will be required to build large-scale argumentation solvers.
Finally, we have advocated CA for contestability, but other KR techniques could provide useful support for some contestability aspects. Specifically, formal verification of AI models could be used to ensure the functional correctness of ADSs, e.g. as in (Kouvaros et al., 2023). Similarly, rigorous logic-based explainability techniques (Narodytska et al., 2019; Darwiche, 2020) could be used to ensure the faithfulness and trustworthiness of explanations.